sir my question is
sin(-660”)tan(1050*)sec(-420”)/cos(225”)cosec(315”)cos(510”)
its answer is given 2/underroot 3
 
how it will come
& “ symbol is representing sign of degree

Grade:12th pass

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
V = \frac{sin(-660)tan(1050)sec(-420)}{cos(225)cosec(315)cos(510)}
V = \frac{sin(-660)tan(1050)sin(315)}{cos(225)cos(-420)cos(510)}
V = \frac{-sin(660)tan(1050)sin(315)}{cos(225)cos(420)cos(510)}
V = \frac{-sin(300)tan(-30)sin(-45)}{-cos(45)cos(60)cos(150)}
V = \frac{sin(300)tan(30)sin(45)}{cos(45)cos(60)cos(150)}
V = \frac{-sin(60)tan(30)sin(45)}{-cos(45)cos(60)cos(30)}
V = \frac{sin(60)tan(30)sin(45)}{cos(45)cos(60)cos(30)}
V = \frac{\frac{\sqrt{3}}{2}.\frac{1}{\sqrt{3}}.\frac{1}{\sqrt{2}}}{\frac{1}{\sqrt{2}}.\frac{1}{2}.\frac{\sqrt{3}}{2}}
V = \frac{2}{\sqrt{3}}